OpenStudy (anonymous):

I don't understand these type of problems.. At all. You don't have to answer them all. Just help me out.. Please? 1. If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4) 2. If f(x) = 1/3x − 2 and g(x) = −2x − 1, find (f • g)(9) 3. If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). 4. If f(x) = 1/2x + 3 and g(x) = −4x − 1, find g[f(4)] 5. If f(x) = x − 4 and g(x) = x − 2, find f[g(−6)] 6. If f(x) = −3x − 3 and g(x) = 2x + 1, find f/g(−2). 7. If r(x) = 3x − 1 and t(x) = −4x − 1, find (r + t)(7) 8. If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)]. Help IS appreciated.

hero (hero):

(f-g)(x) = f(x) - g(x) (f• g)(x) = f(x) • g(x) (f+g)(x) = f(x) + g(x) (f/g)(x) = f(x)/g(x)

hero (hero):

Does that help?

OpenStudy (anonymous):

and about the f(g(x)) substitute the g(x) equation as the f(x) then simplify it

hero (hero):

(f o g)(x) = f(g(x))

OpenStudy (anonymous):

yes

hero (hero):

roundofapplause, if you don't communicate with anyone, we can't help you. No one is going to do all of those for you. Tell us if you get what I posted or not

OpenStudy (anonymous):

I'm sorry, I was eating -_- I understand what you posted, and I appreciate your help. THANK YOU.

hero (hero):

Great :D

OpenStudy (anonymous):

lol